Ketone Inc. produces small engines. For last year's operations, the following data were gathered:

Units produced 110,000
Direct labor 120,000 hours @ $10.00
Actual variable overhead $1,000,000


Ketone Inc. employs a standard costing system. During the year, a variable overhead rate of $5.00 was used. The labor standard requires 1 hour per unit produced. The variable overhead spending and efficiency variances are:

a.$176,000 U and $19,000 U.

b.$400,000 U and $50,000 U.

c.$200,000 U and $40,000 F.

d.$150,000 U and $24,000 F.

e.None of these choices are correct.
Answer #1

Variable overhead spending variance

= Standard voh - Actual voh

= (5*1*110,000) - 1,000,000

= 400,000 U

Voh efficiency variance = (SH - AH) * SR

= (110,000 *1 - 120,000) * 5

= 50,000 U

OPTION B is the answer